There are two rectangular plots of land, A and B. The difference between the length and the width of A is 3 m. The length of A is 8 m more than that of B; the width of A is 5 m less than the width of B and the area of B is 250 m squared more than half the area of A. Find lengths and widths of A and B.
Formula given:
Area of B - 1/2 Area of A = 250

Answer by jaymore(5)

Let the dimensions of plot B be x and y
The dimensions of plot A will be x+8 and y-5
Difference in lengh n width of plot A
(x+8)-(y-5)=3
Solve to get x-y=-10 or x+10=y
Area of Plot A:
(x+8)(y-5)
Expand to get xy-5x+8y-40
Substitute x+10=y in the equation to get
Xsquared+13x+40
Area of Plot B:
xy=x(x+10)=Xsquared+10x
(Xsquared+10x)-1/2(Xsquared+13x+40)=250
Simplyfy to get Xsquared+7x-540=0
Solve quadratic equation
X=20 and y=x+10=20+10=30
Plot A:
Length=28 width=25
Plot B:
Length=20 width=30